Hi I did this problem and I was just wondering if a professional could over look this for me and let me know if it looks okay and if I am missing any steps or anything further! Thank you!!!!! :)
3x+2y=12 for y
3x + 2y = 12
2y = 12 - 3x
y = (12 -3x)/2

Answer by richwmiller(17219)

What you have is fine but I would go a few more steps
y=(12-3x)2
y=6-3x/2
y=6-1.5x